a student opens a mathematics book to two facing pages. the product of the page number is 702. find the page number.
the first page is
the second page is

Found 2 solutions by unlockmath, dlh:
Answer by unlockmath(1688)

Hello,
Let's have x represent the left page and (x+1) be the right page. Now the equation will look like this:
x(x+1)=702
Rewritten as:
x^2+x-702=0
Factor this to get:
(x-26)(x+27)=0
Solve for x:
x=26
x=-27
The sensible answer is 26 so the pages are
26 and 27
Make sense?
RJ
www.math-unlock.com

Answer by dlh(17)

1st page=26
2nd page=27